OpenStudy (anonymous):

Which of the following statements is not true regarding operations with exponents? A)To divide powers with the same base, subtract the exponents. b)To subtract powers with the same base, divide the exponents. c)To multiply powers with the same base, add the exponents. d)To raise a power to a power, multiply the exponents.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Awesome! Thanks~ I'm probs gonna need more a few more questions! I really suck at math. Think you can help?

OpenStudy (anonymous):

x^1 = x x^0 = 1 x^-1 = 1/x x^m*x^n = x^(m+n) x^m/x^n = x^(m-n) (x^m)^n = x^(m*n) (x*y)^n = x^ny^n (x/y)^n = x^n/y^n x^-n = 1/x^n

OpenStudy (anonymous):

Laws of exponents